# arXiv:2305.16494v1[cs.CV] 25 May 2023

### v1

@@ Abstract
Neural networks are known to be susceptible to adversarial samples: small variations of natural examples crafted to deliberately mislead the models. While they can be easily generated using gradient-based techniques in digital and physical scenarios, they often differ greatly from the actual data distribution of natural images, resulting in a trade-off between strength and stealthiness. In this paper, we propose a novel framework dubbed Diffusion-Based Projected Gradient Descent (Diff-PGD) for generating realistic adversarial samples. By exploiting a gradient guided by a diffusion model, Diff-PGD ensures that adversarial samples remain close to the original data distribution while maintaining their effectiveness. Moreover, our framework can be easily customized for specific tasks such as digital attacks, physical-world attacks, and style-based attacks. Compared with existing methods for generating natural-style adversarial samples, our framework enables the separation of optimizing adversarial loss from other surrogate losses (e.g., content/smoothness/style loss), making it more stable and controllable. Finally, we demonstrate that the samples generated using Diff-PGD have better transferability and anti-purification power than traditional gradient-based methods. Code will be released in https://github.com/xavihart/Diff-PGD
@@ 3 Background
Diffusion Models (DMs) [16, 38] have demonstrated superior performance in many tasks such as text-to-image generation [34, 2, 35], video/story generation [18, 17, 32], and 3D generation [25, 33]. Pre-trained DMs provide valuable prior knowledge that can be exploited for adversarial robustness. In this context, several works have proposed defense mechanisms such as adversarial purification [31, 47, 39, 31, 24] or DM-enhanced certified robustness [5, 41, 40].
@@
where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\Pi_{s=1}^{t}\alpha_{s}$. When $\bar{\alpha}_{t}\approx 1$, the distribution of $x_{T}$ becomes an isotropic Gaussian.
@@
The key idea of Stochastic Differential Editing (SDEdit) [29] is to diffuse the original distribution with $x_{K}\sim q(x_{K}\mid x_{0})$ and then run the reverse denoising process $R_{\phi}$ parametrized by $\phi$ in a Diffusion Model to get an edited sample:
@@ 4.1 Diffusion-based Projected Gradient Descent
After $n$ iterations, Diff-PGD generates two outputs ($x^{n}_{0}$ and $x_{n}$) that are slightly different from each other: 1) $x^{n}_{0}$, according to the definition of the loss function, is realistic and adversarial. 2) $x^{n}$ is an optimized perturbed image to which the adversarial gradient is directly applied. It follows the PGD $\ell_{\infty}$ bound, and according to the loss function Equation (5), hard to be purified by SDEdit.
@@ 4.2 Extended Diff-PGD into Region Attack
For the previous gradient-based method, it is easy to modify $\nabla_{x}f_{\theta}(x)$ into $\nabla_{x}f_{\theta}(M\circ x+(1-M)\circ x_{c})$, where $x_{c}=x$ is a copy of the original image with no gradient flows by. It is not surprising the perturbed regions will show a large distribution shift from other regions, making the figure unnatural.
@@ 5 Experiments
Datasets, Models, and Baselines. We use the validation dataset of ImageNet [6] as our dataset to get some statistical results for global attacks and regional attacks. For the style-guided adv-sample generation, we use some cases from [9] but also collect more images and masks by ourselves. We use VGG-19 as the backbone network to compute the style distance. The main target classifier to be attacked across this paper is ResNet-50 [15], and we also use ResNet-101, ResNet18, and WideResNet(WRN) [45] for the transferability test. The diffusion model we adopt is the unconditional diffusion model pre-trained on ImageNet [7] though we use DDIM [37] to respace the original timesteps for faster inference. More details about the experimental settings are in the appendix.
@@
Diff-PGD For Digital Attacks (Q1). We begin with the basic global $\ell_{\infty}$ digital attacks, where we set $\ell_{\infty}=16/255$ for PGD and Diff-PGD, and # of iterations $n=10$ and step size $\eta=2/255$. For Diff-PGD, we use DDIM with timestep $T_{s}=50$ (noted as $\text{DDIM}50$ for simplicity), and $K=3$ for the SDEdit module. Figure 2 shows that adv-samples generated by Diff-PGD ($x^{n}_{0}$) are more stealthy, while PGD-generated samples ($x_{PGD}$) have some patterns that can be easily detected by humans. This contrast is clearer on the perturbation map: $\delta_{PGD}$ contains more high-frequency noise that is weakly related to local patterns of the instance $x$ and $\delta_{0}^{n}$ is smoother and highly locally dependent. Similarly, for regional adversarial attacks, Diff-rPGD can generate adv-samples that can better blend into the unchanged region, as demonstrated in Figure 3, showing higher stealthiness than rPGD.
@@
Diff-PGD For Physical-World Attacks (Q3). In the physical-world attacks, we set the problem as: given an image patch (start point of optimization), and a target object (e.g. attack a backpack), we need to optimize the patch to fool the classifier when attached to the target object. We use an iPhone 8-Plus to take images from the real world and use an HP DeskJet-2752 for color printing. In AdvPatch, only adversarial loss is used to optimize the patch; in AdvCAM, content loss and smooth loss are optimized together with adversarial loss; for Diff-Phys, we only use purified adversarial loss defined in Equation 8, and the SDEdit is set to be DDIM10 with $K=2$ for better purification.
The results are shown in Figure 5, for the untargeted attack, where we use a green apple as a patch to attack a computer mouse. For targetted attack, we use an image of a cat to attack and backpack, with the target-label as “Yorkshire terrier”. We can see that both AdvPatch and AdvCam generate patches with large noise, while Diff-Phys can generate smoother and more realistic patches that can successfully attack the given objects.
@@ 6 Conclusions
In this paper, we propose a novel method to power the generation of adversarial samples with diffusion models. The proposed Diff-PGD method can improve the stealthiness of adv-samples. We further show that Diff-PGD can be easily plugged into global digital attacks, regional digital attacks, customized attacks, and physical-world attacks. We demonstrate through experiments that our methods outperformed the baselines and are effective and stable. The major limitation of our method is that the back-propagation is more expensive than traditional gradient-based methods, however, we believe that the strong prior knowledge of the diffusion model can push forward adversarial attacks & defenses. Finally, while the proposed adversarial attack method could be potentially used by malicious users, it can also enhance future efforts to develop robust defense mechanisms, thereby safeguarding the security of AI systems.
@@
The appendix is organized as follows: in Section A, we provide a summary table of notations to assist readers in better understanding the content. Then we delve into the details of our algorithms in Section B, including how to apply Diff-PGD to style-based attacks and physical world attacks step by step. Then, we show the experimental settings in Section C, including our experimental environments, training settings, and inference speed. Then, we show more complementary results of our experiments in Section D, where we also present the results of a human-evaluation survey. Finally, we show the potential social impacts of our work in Section F.
@@ B.2 Details about Physical World Diff-PGD
Appendix C Implementation Details
@@
Here we use $\epsilon=16/255,\eta=2/255,n=10$ as our major settings (except for the ablation study settings) for both PGD and Diff-PGD. For Diff-PGD, we use DDIM50 with $K_{s}=3$ as our SDEdit setting, we will also show more results with different DDIM time-steps and different $K_{s}$.
@@
For the experiments of the regional attacks, we randomly select a squared mask sized $0.4H\times 0.4W$, where the size of the original image is $H\times W$. For the same image in the ImageNet dataset, we use the same mask for both rPGD and Diff-rPGD. Also, we use $\epsilon=16/255,\eta=2/255,n=10$ for bounded attack with projected gradient descent. And for Diff-rPGD, we use DDIM50 with $K_{s}=2$, similarly, we try more settings in the ablation study section. The repainting strategy of our Diff-rPGD is based on replacement, but more effective strategies are encouraged to be used here.
@@
As mentioned in Section B, the implementation of Diff-PGD for style-based attacks can be divided into two parts. For the first stage, we use style weight $w_{s}=4000$ and $w_{1}$, and adopt Adam optimizer with learning rate $lr=0.01$ for style transfer. For the feature extractor, we adopt VGG-19 as our backbone with the first five convolutional layers as the style extractor $H_{s}$ and the fourth convolutional layer as the content extractor $H_{c}$. For the target image $x$, we use SAM to collect the segmentation mask to define the attack region.
In the second stage, we utilize DDIM10 with a parameter setting of $K_{s}=2$ to achieve better purification. It is beneficial to use a smaller $K_{s}$ as the style-transferred images often contain unwanted noise and art-styled patterns that can be mitigated through this approach.
@@
The background images in the physical world attacks are collected by taking photos of the target object with different camera poses. For the physical adaptor, we set the scale factor to be around $[0.8,1.2]$ of the original patch size, then we set the random brightness to be $[0.5,1.5]$ of the original brightness. We also set a margin for the random positions of the image patch so that the patch will not be too far away from the target object. We use an iPhone 8-Plus to take images from the real world and use an HP DeskJet-2752 for color printing. For Diff-PGD we use DDIM10 and $K_{s}=2$ to allow larger purification.
@@
We show results of two additional good properties of $x^{n}$ and $x^{n}_{0}$ generated by Diff-PGD: transferability and anti-purification power. For transferability, we use $\epsilon=16,\eta=2,n=10$ in the main paper, with DDIM50 with $K_{s}=2$ for Diff-PGD, targeting our network on ResNet-50, we test the adv-samples among ResNet-101, ResNet18, and WRN-50 and WRN-101.
For the test of anti-purification power, we use an SDEdit with DDIM50 with a larger $K_{s}=5$ and test for both PGD and Diff-PGD with $\epsilon=16,\eta=2,n=10$ and DDIM50 with $K_{s}=2$ for Diff-PGD so that our method is not biased.
@@
Appendix D More Results

### v2

# arXiv:2305.16494v2[cs.CV] 11 Jan 2024

@@ Abstract
Neural networks are known to be susceptible to adversarial samples: small variations of natural examples crafted to deliberately mislead the models. While they can be easily generated using gradient-based techniques in digital and physical scenarios, they often differ greatly from the actual data distribution of natural images, resulting in a trade-off between strength and stealthiness. In this paper, we propose a novel framework dubbed Diffusion-Based Projected Gradient Descent (Diff-PGD) for generating realistic adversarial samples. By exploiting a gradient guided by a diffusion model, Diff-PGD ensures that adversarial samples remain close to the original data distribution while maintaining their effectiveness. Moreover, our framework can be easily customized for specific tasks such as digital attacks, physical-world attacks, and style-based attacks. Compared with existing methods for generating natural-style adversarial samples, our framework enables the separation of optimizing adversarial loss from other surrogate losses (e.g., content/smoothness/style loss), making it more stable and controllable. Finally, we demonstrate that the samples generated using Diff-PGD have better transferability and anti-purification power than traditional gradient-based methods. Code is available at https://github.com/xavihart/Diff-PGD
@@ 3 Background
Diffusion Models (DMs) [19, 41] have demonstrated superior performance in many tasks such as text-to-image generation [37, 2, 38], video/story generation [21, 20, 35], 3D generation [28, 36] and neural science [44, 42]. Pre-trained DMs provide valuable prior knowledge that can be exploited for adversarial robustness. In this context, several works have proposed defense mechanisms such as adversarial purification [34, 53, 43, 34, 27] or DM-enhanced certified robustness [6, 46, 45].
@@
where $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\Pi_{s=1}^{t}\alpha_{s}$. When $\bar{\alpha}_{t}\approx 0$, the distribution of $x_{T}$ becomes an isotropic Gaussian.
@@
The key idea of Stochastic Differential Editing (SDEdit) [32] is to diffuse the original distribution with $x_{K}\sim q(x_{K}\mid x)$ and then run the reverse denoising process $R_{\phi}$ parametrized by $\phi$ in a Diffusion Model to get an edited sample:
@@ 4.1 Diffusion-based Projected Gradient Descent
After $n$ iterations, Diff-PGD generates two outputs ($x^{n}_{0}$ and $x^{n}$) that are slightly different from each other: 1) $x^{n}_{0}$, according to the definition of the loss function, is realistic and adversarial. 2) $x^{n}$ is an optimized perturbed image to which the adversarial gradient is directly applied. It follows the PGD $\ell_{\infty}$ bound, and according to the loss function Equation (5), hard to be purified by SDEdit.
@@ 4.2 Extended Diff-PGD into Region Attack
For the previous gradient-based method, it is easy to modify $\nabla_{x}f_{\theta}(x)$ into $\nabla_{x}f_{\theta}(M\circ x+(1-M)\circ x_{c})$, where $x_{c}=x$ is a copy of the original image with no gradient flowing by. It is not surprising the perturbed regions will show a large distribution shift from other regions, making the figure unnatural.
@@ 5 Experiments
Datasets, Models, and Baselines. We use the validation dataset of ImageNet [8] as our dataset to get some statistical results for global attacks and regional attacks. For the style-guided adv-sample generation, we use some cases from [12] but also collect more images and masks by ourselves. We use VGG-19 as the backbone network to compute the style distance. The main target classifier to be attacked across this paper is ResNet-50 [18], and we also use ResNet-101, ResNet18, and WideResNet(WRN) [51] for the transferability test. Besides convolutional-based networks, we also try our Diff-PGD on Vision Transformers like ViT-b [11] and BEiT-l [3], and the results are put in Appendix E. The diffusion model we adopt is the unconditional diffusion model pre-trained on ImageNet [9] though we use DDIM [40] to respace the original timesteps for faster inference. More details about the experimental settings are included in the appendix.
@@
Diff-PGD For Digital Attacks (Q1). We begin with the basic global $\ell_{\infty}$ digital attacks, where we set $\ell_{\infty}=16/255$ for PGD and Diff-PGD, and # of iterations $n=10$ and step size $\eta=2/255$. For Diff-PGD, we use DDIM with timestep $T_{s}=50$ (noted as $\text{DDIM}50$ for simplicity), and $K=3$ for the SDEdit module. Figure 2 shows that adv-samples generated by Diff-PGD ($x^{n}_{0}$) are more stealthy, while PGD-generated samples ($x_{PGD}$) have some patterns that can be easily detected by humans. This contrast is clearer on the perturbation map: $\delta_{PGD}$ contains more high-frequency noise that is weakly related to local patterns of the instance $x$, and $\delta_{0}^{n}$ is smoother and highly locally dependent. Similarly, for regional adversarial attacks, Diff-rPGD can generate adv-samples that can better blend into the unchanged region, as demonstrated in Figure 3, showing higher stealthiness than rPGD.
@@
We also conduct experiments on different norms. We show additional results of the performance of Diff-PGD on $\ell_{2}$-based attacks in Appendix F.
@@
Diff-PGD For Physical-World Attacks (Q3). In the physical-world attacks, we set the problem as: given an image patch (start point of optimization), and a target object (e.g. attack a backpack), we need to optimize the patch to fool the classifier when attached to the target object. We use an iPhone 8-Plus to take images from the real world and use an HP DeskJet-2752 for color printing. In AdvPatch, only adversarial loss is used to optimize the patch; in AdvCAM, content loss and smooth loss are optimized together with adversarial loss; for Diff-Phys, we only use purified adversarial loss defined in Equation (8), and the SDEdit is set to be DDIM10 with $K=2$ for better purification.
The results are shown in Figure 5, for the untargeted attack, where we use a green apple as a patch to attack a computer mouse. For targeted attack, we use an image of a cat to attack a backpack, with the target-label as “Yorkshire terrier”. We can see that both AdvPatch and AdvCam generate patches with large noise, while Diff-Phys can generate smoother and more realistic patches that can successfully attack the given objects.
@@
6 Accelerated Diff-PGD with Gradient Approximation
From Algorithm 1, we can see that we need to calculate the derivative of the model’s output over the input, where the model is composed of chained U-Nets $R_{\phi}$ in the diffusion model followed by a target classifier $f_{\theta}$. Here we focus on the computational bottleneck, which is the derivative of $K$-step SDEdit outputs $x_{0}$ over SDEdit input $x$:
$\frac{\partial x_{0}}{\partial x}=\frac{\partial x_{K}}{\partial x}(\frac{\partial x_{K-1}}{\partial x_{K}}\frac{\partial x_{K-2}}{\partial x_{K-1}}...\frac{\partial x_{1}}{\partial x_{0}})\approx c$ (10)
We approximate the gradient with a constant $c$. We notice that the approximation of this Jacobian has been used in many recent works [36, 49, 7]. We also tried this strategy in Diff-PGD, and we got the accelerated version:
$\frac{\partial L(f(x^{i}_{0}))}{\partial x}\approx c\frac{\partial L(f(x^{i}_{0}))}{\partial x^{n}_{0}}.$ (11)
From this we can see that, we only need to calculate the gradient over $x^{n}_{0}$! We only run the inference of SDEdit to get $x^{n}_{0}$ without saving the gradient (lower GPU memory, faster), making it much cheaper. We found that, for the global attack tasks, the approximated gradient still shows a high success rate but saves $50\%$ of time and $75\%$ of VRAM. More detailed results about the visualization and sucess rate can be found in Appendix C.
@@ 7 Conclusions
In this paper, we propose a novel method to power the generation of adversarial samples with diffusion models. The proposed Diff-PGD method can improve the stealthiness of adv-samples. We further show that Diff-PGD can be easily plugged into global digital attacks, regional digital attacks, customized attacks, and physical-world attacks. We demonstrate through experiments that our methods outperformed the baselines and are effective and stable. The major limitation of our method is that the back-propagation is more expensive than traditional gradient-based methods; however, we believe that the strong prior knowledge of the diffusion model can push forward adversarial attacks & defenses. Finally, while the proposed adversarial attack method could be potentially used by malicious users, it can also enhance future efforts to develop robust defense mechanisms, thereby safeguarding the security of AI systems.
Acknowledgment
The authors would like to thank the anonymous reviewers for useful comments. HX and YC are supported by grants NSF 2008513 and NSF 2206576. AA is supported in part by the Army Research Office under grant number W911NF-21-1-0155 and by the New York University Abu Dhabi (NYUAD) Center for Artificial Intelligence and Robotics, funded by Tamkeen under the NYUAD Research Institute Award CG010. BH is supported by the NSF award CAREER-2048168 and the AFOSR award FA9550-23-1-0732.
@@ Acknowledgment
The appendix is organized as follows: in Section A, we provide a summary table of notations to assist readers in better understanding the content. Then we delve into the details of our algorithms in Section B, including how to apply Diff-PGD to style-based attacks and physical world attacks step by step. Then, we show the experimental settings in Section D, including our experimental environments, training settings, and inference speed. Then, we show more complementary results of our experiments in Section G, where we also present the results of a human-evaluation survey. Finally, we discuss the potential social impacts of our work in Section I.
@@ B.2 Details about Physical World Diff-PGD
Appendix C Accelerated Diff-rPGD
Changing only one line can help us get the accelerated version of generalized Diff-PGD with an approximated gradient (Algorithm 4).
Algorithm 4 Accelerated Diff-rPGD
$x^{0}=x^{0}_{0}=x$, $x_{c}=x$
for $t=0,1,2,\dots,n-1$ do
  for $i=K-1,\dots,0$ do
  end for
  $g=\nabla_{x^{t}_{0}}l[f_{\theta}(M\circ x_{0}^{t}+(1-M)\circ x_{c})]$ $\triangleright$ Compute the approximated gradient
end for
Here, we compared the speed and success rate of accelerated Diff-PGD with the original version; we refer to the accelerated version as v2. The qualitative results of v2 are put in Figure 10, and the quantitative results about speed, VRAM, and success rate are put in Table 3 and Table 4.
Method $K$ $n$ VRAM(G) Speed(sec/sample)
Diff-PGD 2 10 $\sim$18  8
v2 2 10 $\sim$4  4
Diff-PGD 3 10 $\sim$20  10
v2 3 10 $\sim$4  5
Table 3: The Speed and VRAM of Diff-PGD and accelerated Diff-PGD with approximal gradient (v2). From this we can see that the v2 is much cheaper to run than the original Diff-PGD. $K$ and $n$ follows the definition in Algorithm 1.
Method Model $n$ $\epsilon$ SR
v2 ResNet-50 10 16/255 99.6
v2 ResNet-50 10 8/255 98.8
v2 ResNet-50 15 16/255 100.0
v2 ResNet-50 15 8/255 100.0
Table 4: The success rate of accelerated Diff-PGD, from which we can see the success rate (SR) is still really nearly $100\%$, showing that v2 is a cheaper but still effective attack. $\epsilon$ and $n$ follows the definition in Algorithm 1.
Appendix D Implementation Details
@@
Here we use $\epsilon=16/255,\eta=2/255,n=10$ as our major settings (except for the ablation study settings) for both PGD and Diff-PGD. For Diff-PGD, we use DDIM50 with $K_{s}=3$ as our SDEdit setting; we will also show more results with different DDIM time-steps and different $K_{s}$.
@@
For the experiments of the regional attacks, we randomly select a squared mask sized $0.4H\times 0.4W$, where the size of the original image is $H\times W$. For the same image in the ImageNet dataset, we use the same mask for both rPGD and Diff-rPGD. Also, we use $\epsilon=16/255,\eta=2/255,n=10$ for bounded attack with projected gradient descent. And for Diff-rPGD, we use DDIM50 with $K_{s}=2$; similarly, we try more settings in the ablation study section. The repainting strategy of our Diff-rPGD is based on replacement, but more effective strategies are encouraged to be used here.
@@
As mentioned in Section B, the implementation of Diff-PGD for style-based attacks can be divided into two parts. For the first stage, we use style weight $\lambda_{s}=4000$ and adopt Adam optimizer with learning rate $\eta=0.01$ for style transfer. For the feature extractor, we adopt VGG-19 as our backbone with the first five convolutional layers as the style extractor $H_{s}$ and the fourth convolutional layer as the content extractor $H_{c}$. For the target image $x$, we use SAM to collect the segmentation mask to define the attack region.
In the second stage, we utilize DDIM10 with a parameter setting of $K_{s}=2$ to achieve better purification. It is beneficial to use a smaller $K_{s}$, as the style-transferred images often contain unwanted noise and art-styled patterns that can be mitigated through this approach.
@@
The background images in the physical world attacks are collected by taking photos of the target object with different camera poses. For the physical adaptor, we set the scale factor to be around $[0.8,1.2]$ of the original patch size; then we set the random brightness to be $[0.5,1.5]$ of the original brightness. We also set a margin for the random positions of the image patch so that the patch will not be too far away from the target object. We use an iPhone 8-Plus to take images from the real world and use an HP DeskJet-2752 for color printing. For Diff-PGD we use DDIM10 and $K_{s}=2$ to allow larger purification.
@@
We show results of two additional good properties of $x^{n}$ and $x^{n}_{0}$ generated by Diff-PGD: transferability and anti-purification power. For transferability, we use $\epsilon=16/255,\eta=2/255,n=10$ in the main paper, with DDIM50 with $K_{s}=2$ for Diff-PGD. Targeting our network on ResNet-50, we test the adv-samples among ResNet-101, ResNet18, and WRN-50 and WRN-101.
For the test of anti-purification power, we use an SDEdit with DDIM50 with a larger $K_{s}=5$ and test for both PGD and Diff-PGD with $\epsilon=16/255,\eta=2/255,n=10$ and DDIM50 with $K_{s}=2$ for Diff-PGD so that our method is not biased.
@@
Appendix E Results on Vision Transformers
Compared with the Convolutional-based network such as ResNet we used before, Vision Transformers are another branch of strong classifiers that are worth taking into consideration. Here we try a global attack on two famous models on ImageNet, ViT-b (https://huggingface.co/google/vit-base-patch16-224) and BEiT (https://huggingface.co/microsoft/beit-large-patch16-224-pt22k-ft22k).
Figure 9 demonstrates that Diff-PGD can still work well to generate adversarial samples with higher stealthiness. Table 5 shows that Diff-PGD is still effective for vision transformers.
Model $n$ $\epsilon$ SR
ViT-base 10 8/255 100
ViT-base 10 16/255 100
BEiT-large 10 8/255 99.6
BEiT-large 10 16/255 99.6
Table 5: Success rate of Diff-PGD on vision transformers.
Appendix F Results on $\ell_{2}$-based Attacks
We also conduct experiments for $\ell_{2}$-based attacks, we set $n=10$ and $\epsilon(\ell_{2})=16/255$, and we can still get a high success rate as $100\%$ and the generated adversarial samples are shown in Figure 8, which turns out to be also realistic.
Figure 8: Performance of Diff-PGD on $\ell_{2}$ perturbations : all the figures attack ResNet-50 with $\epsilon=16/255$ under $l_{2}$ norm. (Zoom-in for better observation)
Figure 9: Qualitative Results of Diff-PGD on Vision Transformers: the left-half attack ViT-b and the right-half attack BEiT-l. (Zoom-in for better observation)
Figure 10: Qualitative Results of Diff-PGD with Gradient Approximation : here we show six samples, each contains $x,x^{n},x^{n}_{0}$ from left to right. (Zoom-in for better observation)
Figure 11: More Results of Style-based Attacks: we show nine samples from the constructed dataset, each sample contains $x,M,x_{s},x^{n}_{0}$ from left to right.
Appendix G More Complementary Results